OpenStudy (anonymous):

final 6.002x fall 2012

OpenStudy (anonymous):

what u mean by your question??

OpenStudy (anonymous):

I finished all the items, some of them wrong, I only have a chance to get the 60% with the Q6, but I failed two times and I only have one chance. My parametrs in this problem are VO=5V, T=0.5ms RW=1 ohm LW=1mH and RL=1 ohm .

OpenStudy (nurali):

@peruana 0.5 0.5 1.811 0.697

OpenStudy (anonymous):

Hey, you, I finished my exam and I got a 60%, but I have a question, wich were the correct answer to problem Q2( b) The circuit parameters have the following values: VS=4V and R=1kΩ, and the ExpFET parameters are: K=1mA, VT=1V and VN=1V. and VIN =2V My answer was 4.39 and was wrong

OpenStudy (anonymous):

Q2(b): vs-kr(e^((vin-vt)/vn)-1) so; 4-1(e^((2-1)/1)-1)=2.281718

OpenStudy (anonymous):

Thanks nickfrei I saw what was my error. I used VIN=0.1 of part a insted of 2 of part b.

OpenStudy (anonymous):

Q1 The elements in the circuit have the following values:R=3k and rm=12k plz help me for this ..................................

OpenStudy (nurali):

@Neelu Question 2 1. 5 2. 1.56 3. 1.81 4. 2.58 5. -5.43 6. -2

OpenStudy (nurali):

@Neelu if correct please click best response.

OpenStudy (nurali):

@Neela 1. 0.1 2. 0.1 3. 2.9239 4. 1.0753

OpenStudy (nurali):

@Neela if correct please click best response.

OpenStudy (anonymous):

Q1 The elements in the circuit have the following values:R=3k and rm=12k plz help me for this ..................................

OpenStudy (anonymous):

@Neelu Q1. a)ia=2*i1*R/(2*R+rm) and vo=ia*rm b)iA =(-i2)*R/(2*R+rm) and V0=ia*rm c) part a) + part b) d)2R/1000 a) (2*1*3/(2*3000+12000))*12000 === 4 b) ((-1)*3/(2*3000+12000))*12000 === -2 c) 4-2 === 2 d) 2*3 === 6